4. [35 points] The following table contains the ACT scores and the GPA for ten college students. Student GPA ACT 2.821 3.4 24 3.026 3.5 27 3.629 3.025 3.225 3.5 28 2.520 3.7 30 10 (a) [10 points] Estimate the relationship between GPA and ACT; that is, obtain the intercept and slope estimates in the equation (You may use a calculator or a spreadsheet program, but show your work.) (b) [2 points] Does the intercept have a useful interpretation here? Explain. (c) [3 points] How much higher is GPA predicted to be if the ACT score is increased (d) [10 points] Compute the fitted values and residuals for each observation and list (e) [5 points] What is the predicted value of GPA when ACT - 20? Does this seem (f) [5 points] How much of the variation in GPA for these ten students is explained by five points? them in a table. Verify that the residuals sum to zero (approximately) like a reasonable predicted value? by ACT? Explain.
Answer #1

SUMMARY OUTPUT

Regression Statistics

Multiple R

0.9011

R Square

0.8119

Adjusted R Square

0.7884

Standard Error

0.1786

Observations

10

ANOVA

df

SS

MS

F

Significance F

Regression

1

1.10

1.10

34.53

0.00

Residual

8

0.26

0.03

Total

9

1.36

Coefficients

Standard Error

t Stat

P-value

Lower 95%

Upper 95%

Intercept

0.4676

0.4718

0.9912

0.3506

-0.6203

1.5555

ACT

0.1079

0.0184

5.8765

0.0004

0.0656

0.1503

a) GPA = 0.4676+0.1079*ACT

b) As the value of the intercept is low and the p-value is less than 0.05, it is not significant and a chance factor

c) GPA predicted increases by 0.5395=(0.1079*5)

d) Yes, the residuals sum to zero

Student

GPA

ACT

Predicted

Residuals

1

2.8

21

2.7

-0.1

2

3.4

24

3.1

-0.3

3

3

26

3.3

0.3

4

3.5

27

3.4

-0.1

5

3.6

29

3.6

0.0

6

3

25

3.2

0.2

7

3.2

25

3.2

0.0

8

3.5

28

3.5

0.0

9

2.5

20

2.6

0.1

10

3.7

30

3.7

0.0

Sum

0.1

e) GPA =0.4676+0.1079*20 = 2.63. yes as predicted value is close to actual value

f) We need to use R square value to understand the variation in GPA. As the value is 0.8119, 81.19 percent of variation in GPA is explained by ACT.